OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

jigglypuff314 (jigglypuff314):

\[slope = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given coordinates (x1,y1) and (x2,y2) and point-slope form is y - y1 = m(x-x1) so you find the slope and plug in to the format

OpenStudy (anonymous):

so y-0=-3(x-4)

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

what about this question What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3?

jigglypuff314 (jigglypuff314):

well you know how to put it into y - y1 = m(x-x1) right? then you try to get it into ax+by+c = 0 form basically to get from y - y1 = m(x-x1) to standard form, you simplify and move everything to one side :)

OpenStudy (anonymous):

so my first step would be y+2=-3x-4

jigglypuff314 (jigglypuff314):

be careful, there is distributive property so it is y+2=-(3x-4) which equals y+2=-3x+12 parenthesis matter :P

OpenStudy (anonymous):

so no i distribute the -?

jigglypuff314 (jigglypuff314):

oh wait, the slope is -3 so the point-slope equation should be y+2=-3(x-4) which equals y+2=-3x+12 sorry for confusing you there

OpenStudy (anonymous):

so no i subtract 12?

jigglypuff314 (jigglypuff314):

y+2=-3(x-4) y+2=-3(x)+(-3)(-4) -3 * -4 = positive 12

OpenStudy (anonymous):

i ment now i subtract 12 to get it to the other side right?

jigglypuff314 (jigglypuff314):

oh yeah, yes :)

OpenStudy (anonymous):

can i subtract 2 insted?

jigglypuff314 (jigglypuff314):

you can if you want, but by doing that you'll have a bunch of negative values on one side, while if you subtract 12 you it becomes easier on the eyes to look at and comprehend... but it is a preference thing so yes you can :)

OpenStudy (anonymous):

i got 3x+y=10

jigglypuff314 (jigglypuff314):

yes, that is correct :D